A class contains 16 boys and 10 girls of which half of the girls are fat. Find the probability that a student chosen at random is a boy or a fat student.


1
Expert's answer
2021-07-27T10:58:04-0400

Total number of students = 16+10 = 26

Number of fat students = 5 (only girls)

"P = \frac{16}{26} + \frac{5}{26} - \frac{0}{26} \\

= \frac{16+5}{26} \\

= \frac{21}{26}"